OpenStudy (mtalhahassan2):

m+1 divide by 6 + m-3 divide by 3

OpenStudy (aum):

Use the exact same method as in the previous problem.

OpenStudy (mtalhahassan2):

I did bro

OpenStudy (mtalhahassan2):

\[m + 1\div 6 + m-3 \div 3\]

OpenStudy (shinalcantara):

\[\frac{ m + 1 }{ 6 } + \frac{ m-3 }{ 3 }\] first thing to do is to find the LCD

OpenStudy (mtalhahassan2):

6

OpenStudy (shinalcantara):

then you'll have it as \[\frac{ (m + 1) + 2(m - 3) }{ 6 }\]

OpenStudy (shinalcantara):

simplify the numerator

OpenStudy (mtalhahassan2):

how you get 2

OpenStudy (shinalcantara):

\[\frac{ m-3 }{ 3 } \rightarrow \frac{ 2(m - 3) }{ 6 }\] Since you have 6 as the LCD, to find the value for the numerator, all you have to do is divide the LCD with the original denominator in which in that case: 6 is the LCD and 3 is the original denominator.. so you'll have 2 then multiply it to the numerator this is also what happened to the first term \[\frac{ m+1 }{ 6 } \rightarrow \frac{ 1(m+1) }{ 6 }\] only that 6/6 is 1 so won't notice it that much

OpenStudy (mtalhahassan2):

so that be the answer

OpenStudy (mtalhahassan2):

sorry what

OpenStudy (shinalcantara):

notice that when you simplify it to lowest term dividing both the numerator and the denominator by 2 you'll come up with the original expression \[\frac{ 2(m-3) }{ 6 } \rightarrow \frac{ \frac{ 2 }{ 2 }(m-3) }{ \frac{ 6 }{ 2 } } \rightarrow \frac{ m-3 }{3 }\]

OpenStudy (shinalcantara):

\[\frac{ m+1 }{ 6 } + \frac{ m - 3 }{ 3 } = \frac{ (m+1) + 2(m-3) }{ 6 }\] simplifying it further you'll have \[ = \frac{ (m+1) + ((2)m - 2(3)) }{ 6 }\] \[ = \frac{ (m+1) + (2m - 6 }{ 6 } \] \[ = \frac{ m+1 + 2m - 6 }{ 6 } \] \[ = \frac{ 3m - 5 }{ 6 } \]

OpenStudy (mtalhahassan2):

thanks man